# Review on the decoding algorithms for surface codes

### v4

@@ 5.5.3 Neural-network decoders
Neural-network (NN) decoders are a result of applying machine learning (ML) techniques for decoding quantum error correction codes [125]. In order to do so, the authors of the NN decoder proposed reducing the decoding problem to the well-studied ML problem of classification, which generally consist on optimizing the assignment of known labels (generally low-dimensional) to known inputs (generally high-dimensional) with the scope of afterwards using the optimized assignment to label inputs that are unknown. A toy example of a classification problem in ML is to label photos of cats and dogs, i.e. to say which photos depict cats and which dogs [196]. The stage in which the classifier is optimized with known data is named training. Hence, NN decoders need a training stage before being able to decode. In the context of the analogy with machine learning, the authors propose to decompose an error into three multi-qubit Pauli operators as $E=STL$, where $S$ is an element of the stabilizer, $T$ is any fixed Pauli that produces the error syndrome (usually referred as pure error) and $L$ is a logical Pauli operator [125]. Note that this decomposition is standard in the field of quantum error correction [54]. Due to the degenerate nature of QECCs, any recovery operator $E^{\prime}=S^{\prime}TL$ will successfully correct the corrupted codeword implying that the stabilizer element of the recovery operator can be assigned arbitrarily with no impact in the logical error rate of the correcting method. In addition, the authors discuss that the pure error can be produced using a parallel table look-up since each element of the error syndrome depends on a unique pure error independently of the others [125]. The algorithms performing such two assignments is named as the simple decoder. Hence, the NN decoder is based on the fact that since the rotated planar code encodes one logical qubit, the logical operators associated are $\mathrm{\hat{I},\hat{X},\hat{Y},\hat{Z}}$ (the hat refers to the fact that those operators are defined over the logical qubit) and, therefore, decoding can be done with as a classification problem with those four labels.
@@
The way in which the classification problem was dealt with in [125] was by using feed-forward neural networks. The neural network in question is defined by a graph consisted of many artificial neuron layers which are connected among them. The neurons are defined by a so called-activation function which depends on an array of weights, $\bar{w}$, of length $m$ and a bias $b$. Since the NN is of the feed forward type, the neurons calculate the nonlinear activation function ($\bar{x}$ is the input and * refers to the inner product), $\bar{y}=(1+\mathrm{exp}(-(\bar{w}*\bar{x}+b)))^{-1}$, and pass it to the subsequent layer until the output layer is reached, where the result of the activation function is rounded to $\{0,1\}$ so that the output of the algorithm is obtained [125]. The layers between the input and output layers are named as hidden layers in the terminology of ML. As explained before, a neural network requires a training stage before it can operate over a set of unknown data. Therefore, for the NN decoder, a training stage where the weights and biases of the neurons are trained is needed. The authors of [125] select the average cross-entropy as the cost function to optimize the activation function and produce a training set directly sampling errors at an error probability where the MWPM decoder has a $25\%$ logical error rate. The cost-function was minimized by stochastic gradient descent, while the number of elements of the training set was limited to at most $10^{6}$ samples.
@@
Figure 24: Graphical depiction of the feed-forward neural network used for the NN decoder. The example consists of an input layer (i), a hidden layer (h) and an output layer (o). The inputs represent the syndrome data, while the outputs refer to the estimated logical error. Each of the neurons in every layer compute the activation function as a function of their received information and pass the information to the next layer. Figure taken from [125].
@@ 5.5.4 MaxSAT decoder
The MaxSAT decoder has been recently proposed as an efficient algorithm to decode color codes [126]. This decoder is based on the analogy between the decoding problem of the color code and the LightsOut puzzle when the error model in consideration is the bit-flip noise. The LightsOut puzzle refers to a problem where a lattice whose faces are associated with switches and lights can be either on or off. In such lattice, toggling a switch on the lattice implies that all its neighbouring lights change their previous state, i.e. if they were on they are turned of and vice versa. The LightsOut problem then tries to find out a sequence of switch actions so that all lights are turned off. The puzzle has two important properties: toggling a switch twice is the same as not doing anything and the state of a light only depends on how often its neighbour switch have been toggled (independence on the order). The authors argue that this problem is similar to the decoding process of a color code by doing the following analogy:
@@
where $f\in F$ refer to the lights or faces, the $\mathrm{switch}_{v}$ are Boolean variables representing if a switch is toggled, $\mathcal{F}_{\mathrm{switches}}(f)$ is a discrete function that takes a face $f$ and returns the set of switches surrounding it; and $\mathcal{F}_{\mathrm{init}}$ is a Boolean function that describes the initial configuration of the system (associated with the measured syndrome). $\bigoplus$ denotes the exclusive-or (XOR) Boolean operation. In equation (21), the constraints in the first line represent the satisfiability problem, and the second are the soft constraints so that such problem is maximum. Hence, color code decoding can be done by solving such MaxSAT problem [126]. The authors discuss many MaxSAT solvers for decoding the color code.
@@ 5.6 Software Packages
In this section an overview regarding the existing software packages for surface code decoding simulation will be covered. Fortunately, the scientific community has provided a large number of open source repositories for performing simulations of several decoders of the surface code and other topological codes and, thus, we will cover several of them (the most popular and used ones) as a reference for people on the community of quantum error correction. We provide a graphical overview in FIG. 26.
@@
Figure 26: Sketch representing the available software packages for each mentioned decoder. Each edge connects each decoder (green circle), with the software packages which support it (pink circles). The decoders mentioned in the figure are the Cellular Automata decoder (CA), the MaxSAT decoder, the Union Find decoder (UF), the Minimum Weight Perfect Matching decoder (MWPM), the Tensor Network decoder (TN) the Neural Network decoder (NN), the Belief Propagation + Order Statistics Decoder (BPOSD) and the Renormalized Group decoder (RG).
@@
Whenever considering numerical simulations of a surface code, it is important to be able to generate the samples of the noise following the distribution of interest for a certain scenario so that the decoder operates over such noise model. In this sense, generating samples for data qubit noise is pretty straightforward by considering the specific distribution of the Pauli channel for each of the qubits (See section 4). Nevertheless, whenever circuit-level noise for faulty check measuments is considered, sampling the noise is not a trivial task [136]. Therefore, the Stim repository by Craig Gidney was developed for sampling circuit-level noise for the sake of obtaining samples of noise that accurately resemble the errors that a surface code experiences over such scenario [136]. Stim samples circuit noise to all qubits within the code, including measurement qubits, at a fast speed and is the simulator of choice when considering stabilizing circuit noise. This repository can analyze a $d=100$ surface code circuits in approximately 15 seconds for sampling circuit shots at a rate of 1000 samples per second.
@@
Regarding software implementations of the minimum weight perfect matching decoder there are several open source repositories which successfully simulate realizations of the decoder at a sufficient speed. The most popular and quickest one considering serial computation is Oscar Higgot’s the Pymatching repository [57, 120], the latest version of which, implements the Sparse Blossom implementation of the MWPM decoder. The Fusion Blossom algorithm [119] is also available at the GitHub repository by Yue Wu [118] which, as explained before, albeit having a worse serial performance than Pymatching, allows parallel decoding improving the complexity at each additional node and eventually surpassing the performance of Pymatching. Additionally, the QECSIM repository by David Tuckett [180] also stands as a fair simulator for the minimum weight perfect matching which also allows other decoding methods. Another recent software package that incorporates Pymatching and Fusion Blossom (by calling those repositories) was posted in the Plaquette repository by the QC Design team [216]. Interestingly, such repository incorporates Stim too, allowing the combination of all those repositories in a single one.
@@
An implementation of the UF decoder can be found in the Qsurface [217]. The package allows the simulation of the standard union-find decoder [109] as well as a modification of it by the authors of Qsurface which they name as Union-find Node-Suspension decoder. The authors claim that an improved threshold can be obtained by using such modified decoder while maintaining complexity. Moreover, such package includes an implementation of the MWPM algorithm. The previously discussed Plaquette package [216] also includes an implementation of the UF decoder by Delfosse and Nickerson. The inclusion of this decoder is specially interesting for the Plaquette package since the circuit-level noise obtained from Stim can be directly used for testing the UF decoder.
@@
The BPOSD algorithm was originally thought as a general decoder for QLDPC codes [46]. However, a surface code can be seen as a sparse code as the distance of the code increases (due to the fact that the weight of the stabilizer is constant independently of the size). In this sense, the BP+OSD: A decoder for quantum LDPC codes package by Joschka Roffe [218] implements the BPOSD decoder for general QLDPC codes. For simulating surface codes, the parity check matrix [30] of such code should be used whenever defining the QLDPC code to be decoded by the BPOSD implementation.
@@
The aforementioned QECSIM repository [180] does also include an implementation of the tensor network decoder. The package include the possibility of tuning the truncation parameter that is used for the TN decoder. Another implementation of the tensor network decoder can be found in the SweepContractor repository by Cristopher Chubb [219] which is based in the sweep line algorithm of [171].
@@
In the case of cellular automaton decoders, the repository Sweep-Decoder-Boundaries by Michael Vasmer [220] implements the version of the decoder using the sweep rule as presented in [185, 186]. As discussed in the papers, the implementation of the decoder admits simulations using circuit-level noise. Additionally, the repository LocalToricDecoder by Kasper Duivenvoorden implements the version of the cellular automaton using Toom’s rule as proposed in [183, 184]. We have been unable to find software packages implementing the other cellular automaton decoders discussed.
@@
Concerning the RG decoder, the QTop repository by Jacob Marks [221] presents an implementation of it. However, it is noteworthy to say that the author stated in the repository that such software package is not being actively maintained and, thus, it may be unreliable. Anyway, the code is available there for anyone that may be interested in testing it or using it to code its own version of RG decoders.
@@
With respect to neural network decoders, the repository Neural Network Decoders for Quantum Error Correcting Codes by Stefan Krastanov [222] implements the neural network decoder proposed in [204] for the toric code. Moreover, Pooya Ronagh developed a the repository Deep Neural Decoders for Fault-Tolerant Quantum Error Correction [223] for the deep neural network decoder proposed in [197] for decoding rotated planar codes. Interestingly, the authors of both repositories programmed the decoders so that other neural networks than the ones in [204, 197] can be incorporated. Hence, those repositories provide freedom to integrate other neural network decoders of the literature. Finally, the repository neural_network_decoder by Paul Baireuther [224]contains the implementation of the neural network decoder for the color code in [212].
@@
To finish with this section, there is a repository with the implementation of the MaxSAT decoder presented in [126]. The repository, developed at the Technical University of Munich, is named QECC: An MQT tool for Quantum Error Correcting Codes written in C++ [225] and is intended to be a more general tool for QEC than just the MaxSAT decoder for color codes. At the time of writing, the authors of the repository state that the project is still in early development and, thus, will contain other QEC software tools. The main ideas used for the software package were presented in [226]. Either way, at the moment of writing, the MaxSAT decoder and a decoder for QLDPCs based on the UF decoder [140] are present in the software package. Note that since surface codes can be seen as QLDPC codes, the UF decoder implemented in [225] can in principle be used to decode such families of codes.
@@ 6 Discussion
As presented in the main text, the main decoders for the rotated planar code family are the MWPM, UF, BPOSD and TN decoders. FIG. 27 is an illustrative representation of the accuracy versus runtime trade-off discussed through the text for those decoders. At the time of writing, it seems that the MWPM is an strong candidate due to its high threshold and the fact that it always returns the $X$ and $Z$-errors of minimum weight matching the observed syndrome [^28] Note that this is not the same as returning the overall minimum weight errors since $X$ and $Z$ errors are evaluated independently and, as such, $Y$ errors are considered as events of double weight as opposed to Pauli errors of weight 1.. In addition, the MWPM decoder presents a considerable worst-case complexity when compared to other decoding methods, but recent implementations of the algorithm have proven to present an almost linear expected complexity [57, 119], making this method even more powerful yet. On the other hand, UF also stands as a fair contestant due to its decoding speed which comes from a linear worst-case complexity. Nevertheless, as discussed before, the MWPM implementations with an average complexity lower than $\mathcal{O}(n^{2})$ [57], even reaching linearity given that enough nodes for parallelization are available [119], make the candidacy of the UF decoder to significantly lose strength. Anyway, it is noteworthy to state that many of those fast MWPM implementations were constructed based on many ideas introduced by UF decoding. The BPOSD-0 decoder lays in a middle ground between the MWPM and UF decoders whenever the rotated planar code is considered. As discussed before, increasing the search depth of the OSD algoirthm has a considerable impact on the worst-case complexity of the decoder and, thus, the benefits of this decoder will probably be lost due to the fact that the threshold of the code would not increase as much. In this sense, it seems that the MWPM decoder is a better candidate for the specific instance of surface codes. Nonetheless, the BPOSD algorithm was proposed as a general decoder for QLDPC codes [46] and, therefore, it is the best candidate for decoding such family of codes which is being studied thoroughly at the time of writing. As seen in the earlier section, its implementation allows the circumnavigation of split beliefs in conventional belief propagation decoding, which have posed a great hurdle to conventional BP decoding for QLDPC codes [150, 46]. The TN decoder achieves the highest code threshold for the rotated planar code, albeit at the expense of the largest worst-case computational complexity. Both the benefit and drawback of this algorithm is a result of the fact that it is an effort to solve the maximum-likelihood (ML) problem explicitly, and this is why it is usually referred as the ML decoder [^29] Note that in the terminology of Section 2 the TN actually solves the DQMLD problem, but the literature on the TN decoder refers to this just as ML.. In this sense, the TN decoder is an almost brute force approach where such brute force search is limited by the bond dimension, $\chi$, which is the approximation parameter that is needed for this method to be actually implementable for codes of considerable size. Hence, the bond dimension is the parameter that will be in charge of the accuracy and speed of the algorithm. Due to the enormous growth in bond dimension in the surface code tensor network contraction, the tensor network is not really considered for real-time decoding but is studied cautiously as an effective degenerate quantum maximum likelihood decoder. Also this decoder presents the possibility of the experimental implementation of codes beyond break-even with poorer hardware capabilities due to its high threshold. The conclusion to this is that the choice of the decoder for a future real-time implementation of surface codes remains as an open question. This quandary is even more complex taking into account the existence of other decoding approaches such as the CA, RG, NN or MaxSAT decoders discussed in this tutotrial. Each of those may provide interesting features that may be beneficial for specific families of codes or even adaptations of them may be interesting for the context of rotated planar codes. For example, machine learning base decoders are gaining popularity on the recent times, probably due to the huge interest in machine learning in general.
@@
In this line, recent breakthroughs have been achieved in the experimental implementations of surface codes over superconducting qubit platforms [63, 64]. Both of the experimental implementations were based on the rotated planar code discussed throughout this tutorial. In [63], Andreas Wallraff’s group at ETH Zurich led an experiment based on the rotated planar code of distance three by means of a processor consisted of 17 transmon qubits. Krinner et al. used the obtained data after multiple measurement rounds in order to run the MWPM decoder for estimating the errors corrupting the quantum information. In [64], the Google Quantum AI realized an experiment based on the rotated planar codes of distances 3 and 5 by means of their expanded Sycamore device consisted of 72 transmon qubits. The authors actually run experiments for the distance-5 rotated planar codes since the actual data for the distance-3 code can be obtained from such code. Acharya et al. used the obatined data to decode the errors by means of the TN and belief-matching decoders [^30] Recall that the belief-matching decoder combines the output of a BP decoder in order to reweight the edges of the graph in the MWPM problem [87].. In this way, the authors proved that the logical error rate obtained by the rotated planar code improved as the distance of the code increased for both decoders. This result is really important since the code was operating in a sub-threshold noise level and it represents an experimental proof that the performance of the codes can be increased by increasing the distance (recall the threshold theorem). Importantly, both of the experiments performed the execution of the decoding algorithm by dealing with the measured data in a post-processing stage, i.e. no real-time decoding was actually performed. However, both of these experiments represent the state-of-the-art of experimental implementations of surface codes and they are an effort to progress towards the ultimate goal of fault-tolerant quantum computes.

### v5

@@ 5.5.3 Neural-network decoders
Neural-network (NN) decoders are a result of applying machine learning (ML) techniques for decoding quantum error correction codes [101]. In order to do so, the authors of the NN decoder proposed reducing the decoding problem to the well-studied ML problem of classification, which generally consists on optimizing the assignment of known labels (generally low-dimensional) to known inputs (generally high-dimensional) with the scope of afterwards using the optimized assignment to label inputs that are unknown. A toy example of a classification problem in ML is to label photos of cats and dogs, i.e. to say which photos depict cats and which dogs [161]. The stage in which the classifier is optimized with known data is named training. Hence, NN decoders need a training stage before being able to decode. In the context of the analogy with machine learning, the authors propose to decompose an error into three multi-qubit Pauli operators as
$E=STL,$ (41)
where $S$ is an element of the stabilizer, $T$ is any fixed Pauli that produces the error syndrome (usually referred as pure error , and it is the part of the error that produces the syndrome [41]) and $L$ is a logical Pauli operator [101]. Note that this decomposition is standard in the field of quantum error correction [41]. FIG. 29 shows an example of the decomposition in equation (41), note that the triggered syndrome is associated to the $T$ component of the error. Due to the degenerate nature of QECCs (Recall that errors up to an stabilizer act equivalently on the logical qubits), any recovery operator $E^{\prime}=S^{\prime}TL$ will successfully correct the corrupted codeword implying that the stabilizer element of the recovery operator can be assigned arbitrarily with no impact in the logical error rate of the correcting method. In addition, the authors discuss that the pure error can be produced using a parallel table look-up since each element of the error syndrome depends on a unique pure error independently of the others [101]  [^34] This has been previously seen in the TN decoder, where $E_{rec}$ solves the syndrome by matching every non-trivial syndrome element to the nearest virtual check.. In this sense, the authors are able to remove the pure error part. The algorithms performing such two assignments (selection of $S$ and $T$ components) are named as the simple decoder. Hence, the NN decoder is based on the fact that since the rotated planar code encodes one logical qubit, the possible logical operators associated are $\mathrm{\hat{I},\hat{X},\hat{Y},\hat{Z}}$ (the hat refers to the fact that those operators are defined over the logical qubit) and, therefore, decoding can be done by means of a classification problem with those four labels.
The way in which the classification problem was dealt with in [101] was by using feed-forward neural networks. Neural networks are machine learning models that aim to mimic the neuron structures of biological brains. In this sense, neural networks are defined by a graph consisted of many artificial neuron layers (nodes) which are connected among them (edges). The neurons are defined by a so called-activation function (that defines the interactions between the artificial networks) which depends on an array of weights, $\bar{w}$, of length $m$ and a bias $b$. Since the NN is of the feed-forward type, the neurons calculate the nonlinear activation function
$\bar{y}=(1+\mathrm{exp}(-(\bar{w}*\bar{x}+b)))^{-1},$ (42)
@@
where $\bar{x}$ is the input (the measured syndrome in the first layer) and * refers to the inner product, and pass it to the subsequent layer until the output layer is reached [^35] Since the considered NN was feed-forward the data exchanged among layers is done in just one direction, from the input to the output. There are instances of NNs that consider both directions for the information flow. Those are referred to as recurrent neural networks., where the result of the activation function is rounded to $\{0,1\}$[101]. The combinations of the outputs define which logical error is estimated and combined with the results of the simple decoder for estimating the complete recovery operator. Note that the estimated error will match the syndrome, but it is not guaranteed to be of minimum-weight due to the heuristic nature of the neural network method. The layers between the input and output layers are named as hidden layers in the terminology of ML. As explained before, a neural network requires a training stage before it can operate over a set of unknown data. Therefore, for the NN decoder, a training stage where the weights and biases of the neurons are trained is needed. The authors of [101] select the average cross-entropy as the cost function to optimize the activation function and produce a training set directly sampling errors at an error probability where the MWPM decoder has a $25\%$ logical error rate. The cost-function was minimized by stochastic gradient descent, while the number of elements of the training set was limited to at most $10^{6}$ samples.
@@
Figure 30: Graphical depiction of the feed-forward neural network used for the NN decoder. The example consists of an input layer (i), a hidden layer (h) and an output layer (o). The inputs represent the syndrome data, while the outputs refer to the estimated logical error. Each of the neurons in every layer compute the activation function (for this example it is given in equation (42)) as a function of their received information and pass the information to the next layer. Figure taken from [101].
@@ 5.5.4 MaxSAT decoder
The MaxSAT decoder has been recently proposed as an efficient algorithm to decode color codes [102]. This decoder is based on the analogy between the decoding problem of the color code and the LightsOut puzzle when the error model in consideration is the bit-flip noise. The LightsOut puzzle refers to a problem where a lattice whose vertices are associated with switches and faces with lights that can be either on or off. In such lattice, toggling a switch on the lattice implies that all its neighbouring lights change their previous state, i.e. if they were on, they are turned off and vice versa. The LightsOut problem then tries to find out a sequence of switch actions so that all lights are turned off. The puzzle has two important properties: toggling a switch twice is the same as not doing anything and the state of a light only depends on how often its neighbour switch have been toggled (independence on the order). The authors argue that this problem is similar to the decoding process of a color code by doing the following analogy:
@@
where $f\in F$ refer to the lights or faces, the $\mathrm{switch}_{v}$ are Boolean variables representing if a switch is toggled, $\mathcal{F}_{\mathrm{switches}}(f)$ is a discrete function that takes a face $f$ and returns the set of switches surrounding it; and $\mathcal{F}_{\mathrm{init}}$ is a Boolean function that describes the initial configuration of the system (associated with the measured syndrome). $\bigoplus$ denotes the exclusive-or (XOR) Boolean operation. In equation (43), the constraints in the first line represent the satisfiability problem, and the second are the soft constraints so that such problem is maximum. Thus, as a result of the analogy presented before, solving such problem leads to an error of minimum weight (from the soft constraints) that matches the syndrome (from the satisfiability problem). Hence, color code decoding can be done by solving such MaxSAT problem [102]. The authors discuss many MaxSAT solvers for decoding the color code.
@@ 5.6 Software Packages
In this section an overview regarding the existing software packages for surface code decoding simulation will be covered. Fortunately, the scientific community has provided a large number of open source repositories for performing simulations of several decoders of surface codes and other topological codes and, thus, we will cover several of them (the most popular and used ones) as a reference for people on the community of quantum error correction. We provide a graphical overview in FIG. 32.
@@
Figure 32: Sketch representing the available software packages for each mentioned decoder. Each edge connects each decoder (green circle), with the software packages which support it (pink circles). The decoders mentioned in the figure are the Cellular Automata decoder (CA), the MaxSAT decoder, the Union Find decoder (UF), the Minimum Weight Perfect Matching decoder (MWPM), the Tensor Network decoder (TN) the Neural Network decoder (NN), the Belief Propagation + Order Statistics Decoder (BPOSD) and the Renormalization Group decoder (RG).
@@
Whenever considering numerical simulations of a surface code, it is important to be able to generate the samples of the noise following the distribution of interest for a certain scenario so that the decoder operates over such noise model. In this sense, generating samples for data qubit noise is pretty straightforward by considering the specific distribution of the Pauli channel for each of the qubits (See section 4). Nevertheless, whenever circuit-level noise for faulty check measurements is considered, sampling the noise is not a trivial task [111]. Therefore, the Stim repository by Craig Gidney was developed for sampling circuit-level noise for the sake of obtaining samples of noise that accurately resemble the errors that a surface code experiences over such scenario [111]. Stim samples circuit noise to all qubits within the code, including measurement qubits, at a fast speed and is the simulator of choice when considering stabilizing circuit noise. This repository can analyze a $d=100$ surface code circuits in approximately 15 seconds for sampling circuit shots at a rate of 1000 samples per second.
@@
Regarding software implementations of the MWPM decoder there are several open source repositories which successfully simulate realizations of the decoder at a sufficient speed. The most popular and quickest one considering serial computation is Oscar Higgot’s the Pymatching repository [44, 97], the latest version of which, uses the sparse blossom implementation of the MWPM decoder. The Fusion Blossom algorithm [96] is also available at the GitHub repository [^36] https://github.com/yuewuo/fusion-blossom by Yue Wu which, as explained before, albeit having a worse serial performance than Pymatching, allows parallel decoding improving the complexity at each additional node and eventually surpassing the performance of Pymatching. Additionally, the QECSIM repository [^37] https://github.com/qecsim/qecsim by David Tuckett also stands as a fair simulator for the minimum weight perfect matching which also allows other decoding methods. Another recent software package that incorporates Pymatching and Fusion Blossom (by calling those repositories) was posted in the Plaquette repository [^38] https://github.com/qc-design/plaquette by the QC Design team. Interestingly, such repository incorporates Stim too, allowing the combination of all those repositories in a single one.
@@
An implementation of the UF decoder can be found in the Qsurface [^39] https://github.com/watermarkhu/qsurface. The package allows the simulation of the standard union-find decoder [89] as well as a modification of it by the authors of Qsurface which they name as Union-find Node-Suspension decoder. The authors claim that an improved threshold can be obtained by using such modified decoder while maintaining complexity. Moreover, such package includes an implementation of the MWPM algorithm. The previously discussed Plaquette package also includes an implementation of the UF decoder by Delfosse and Nickerson. The inclusion of this decoder is specially interesting for the Plaquette package since the circuit-level noise obtained from Stim can be directly used for testing the UF decoder.
@@
The BPOSD algorithm was originally thought as a general decoder for QLDPC codes [37]. However, a surface code can be seen as a sparse code as the distance of the code increases (due to the fact that the weight of the stabilizer is constant independently of the size). In this sense, the LDPC package [^40] https://github.com/quantumgizmos/ldpc by Joschka Roffe implements the BPOSD decoder for general QLDPC codes. For simulating surface codes, the parity check matrix [39] of such code should be used whenever defining the QLDPC code to be decoded by the BPOSD implementation. This package is intended as a module for building and benchmarking LDPC codes. At this moment, the package includes the BPOSD decoder, but a new version of it is expected with implementations of UF and modified BP decoders, among others. We do not include the dependencies to those other decoders in FIG. 32 since the release is not out yet.
@@
The aforementioned QECSIM repository [143] does also include an implementation of the tensor network decoder. The package include the possibility of tuning the truncation parameter that is used for the TN decoder. Another implementation of the tensor network decoder can be found in the SweepContractor repository [^41] https://github.com/chubbc/SweepContractor.jl by Cristopher Chubb which is based in the sweep line algorithm of [142].
@@
In the case of cellular automaton decoders, the repository [^42] https://github.com/MikeVasmer/Sweep-Decoder-Boundaries Sweep-Decoder-Boundaries by Michael Vasmer implements the version of the decoder using the sweep rule as presented in [150, 151]. As discussed in the papers, the implementation of the decoder admits simulations using circuit-level noise. Additionally, the repository LocalToricDecoder by Kasper Duivenvoorden implements the version of the cellular automaton using Toom’s rule as proposed in [148, 149]. We have been unable to find software packages implementing the other cellular automaton decoders discussed.
@@
Concerning the RG decoder, the QTop repository [^43] https://github.com/jacobmarks/QTop by Jacob Marks presents an implementation of it. However, it is noteworthy to say that the author stated in the repository that such software package is not being actively maintained and, thus, it may be unreliable. In any case, the code is available there for anyone that may be interested in testing it or using it to code its own version of RG decoders.
@@
With respect to neural network decoders, the repository [^44] https://github.com/Krastanov/neural-decoder Neural Network Decoders for Quantum Error Correcting Codes by Stefan Krastanov implements the neural network decoder proposed in [169] for the toric code. Moreover, Pooya Ronagh developed another repository [^45] https://github.com/pooya-git/DeepNeuralDecoder Deep Neural Decoders for Fault-Tolerant Quantum Error Correction for the deep neural network decoder proposed in [162] for decoding rotated planar codes. Interestingly, the authors of both repositories programmed the decoders so that other neural networks than the ones in [162, 169] could be incorporated. Hence, those repositories provide freedom to integrate other neural network decoders of the literature. Finally, the repository [^46] https://github.com/baireuther/neural_network_decoder neural_network_decoder by Paul Baireuther contains the implementation of the neural network decoder for the color code in [177].
@@
To finish with this section, there is a repository with the implementation of the MaxSAT decoder presented in [102]. The repository [^47] https://github.com/cda-tum/mqt-qecc, developed at the Technical University of Munich, is named QECC: An MQT tool for Quantum Error Correcting Codes written in C++ and is intended to be a more general tool for QEC than just the MaxSAT decoder for color codes. At the time of writing, the authors of the repository state that the project is still in early development and, thus, will contain other QEC software tools. The main ideas used for the software package were presented in [181]. Either way, at the moment of writing, the MaxSAT decoder and a decoder for QLDPCs based on the UF decoder [115] are present in the software package.
@@ 6 Discussion
As presented in the main text, the main decoders for the rotated planar code family are the MWPM, UF, BPOSD and TN decoders. FIG. 33 is an illustrative representation of the accuracy versus runtime trade-off discussed through the text for those decoders. At the time of writing, it seems that the MWPM is an strong candidate due to its high threshold and the fact that it always returns the $X$ and $Z$-errors of minimum weight matching the observed syndrome [^48] Note that this is not the same as returning the overall minimum weight errors since $X$ and $Z$-errors are evaluated independently and, as such, $Y$-errors are considered as events of double weight as opposed to Pauli errors of weight 1.. In addition, the MWPM decoder presents a considerable worst-case complexity when compared to other decoding methods, but recent implementations of the algorithm have proven to present an almost linear expected complexity [44, 96], making this method even more powerful yet. On the other hand, UF also stands as a fair contestant due to its decoding speed which comes from a linear worst-case complexity. Nevertheless, as discussed before, the MWPM implementations with an average complexity lower than $\mathcal{O}(n^{2})$ [44], even reaching linearity given that enough nodes for parallelization are available [96], make the candidacy of the UF decoder to significantly lose strength. In any case, it is noteworthy to state that many of those fast MWPM implementations were constructed based on many ideas introduced by UF decoding. The BPOSD-0 decoder lays in a middle ground between the MWPM and UF decoders whenever the rotated planar code is considered. As discussed before, increasing the search depth of the OSD algoirthm has a considerable impact on the worst-case complexity of the decoder and, thus, the benefits of this decoder will probably be lost due to the fact that the threshold of the code would not increase as much. In this sense, it seems that the MWPM decoder is a better candidate for the specific instance of surface codes. Nonetheless, the BPOSD algorithm was proposed as a general decoder for QLDPC codes [37] and, therefore, it is the best candidate for decoding such family of codes which is being studied thoroughly at the time of writing. As seen in the earlier section, its implementation allows the circumnavigation of split beliefs in conventional belief propagation decoding, which have posed a great hurdle to conventional BP decoding for QLDPC codes [37, 125]. The TN decoder achieves the highest code threshold for the rotated planar code, albeit at the expense of the largest worst-case computational complexity. Both the benefit and drawback of this algorithm is a result of the fact that it is an effort to solve the maximum-likelihood (ML) problem explicitly, and this is why it is usually referred as the ML decoder [^49] Note that in the terminology of Section 2 the TN actually solves the DQMLD problem, but the literature on the TN decoder refers to this just as ML.. In this sense, the TN decoder is an almost brute force approach where such brute force search is limited by the bond dimension, $\chi$, which is the approximation parameter that is needed for this method to be actually implementable for codes of considerable size. Hence, the bond dimension is the parameter that will be in charge of the accuracy and speed of the algorithm. Due to the enormous growth in bond dimension in the surface code tensor network contraction, the tensor network is not really considered for real-time decoding but is studied cautiously as an effective degenerate quantum maximum likelihood decoder. Also this decoder presents the possibility of the experimental implementation of codes beyond break-even with poorer hardware capabilities due to its high threshold. The conclusion to this is that the choice of the decoder for a future real-time implementation of surface codes remains as an open question. This quandary is even more complex taking into account the existence of other decoding approaches such as the CA, RG, NN or MaxSAT decoders discussed in this tutorial. Each of those may provide interesting features that may be beneficial for specific families of codes or even adaptations of them may be interesting for the context of rotated planar codes. For example, machine learning based decoders are gaining popularity on the recent times, probably due to the huge interest in machine learning in general.
@@
In this line, recent breakthroughs have been achieved in the experimental implementations of surface codes over superconducting qubit and neutral atom platforms [50, 51, 52]. Those experimental implementations were based on the rotated planar code discussed throughout this tutorial. In [50], Andreas Wallraff’s group at ETH Zurich led an experiment based on the rotated planar code of distance-3 by means of a processor consisted of 17 transmon qubits. Krinner et al. used the obtained data after multiple measurement rounds in order to run the MWPM decoder for estimating the errors corrupting the quantum information. In [51], the Google Quantum AI realized an experiment based on the rotated planar codes of distances 3 and 5 by means of their expanded Sycamore device consisted of 72 transmon qubits. The authors actually run experiments for the distance-5 rotated planar codes since the actual data for the distance-3 code can be obtained from such code. Acharya et al. used the obatined data to decode the errors by means of the TN and belief-matching decoders [^50] Recall that the belief-matching decoder combines the output of a BP decoder in order to reweight the edges of the graph in the MWPM problem [65].. In this way, the authors proved that the logical error rate obtained by the rotated planar code improved as the distance of the code increased for both decoders. This result is really important since the code was operating in a sub-threshold noise level and it represents an experimental proof that the performance of the codes can be increased by increasing the distance (recall the threshold theorem). Furthermore, Mikhail Lukin’s group at Harvard, with collaborators of QuEra computing, conducted QEC experiments in a neutral atom processor [52]. Among other relevant QEC experiments, Bluvstein et al. implemented a transversal CNOT gate between the logical qubits of two rotated planar codes for distances $d=\{3,5,7\}$. Importantly, the authors used the transversal CNOT to do a Bell-pair experiment and showed that the experiment error decreases with code distance. It is important to say that the authors discussed the application of a correlated decoder, that considers decoding both logical qubits jointly rather than independently, in order to obtain such result. Two correlated decoding algorithms are proposed, based on belief-find and TN decoders. The experiments used a single round of syndrome measurements, leaving multiple-round extraction as future work. Importantly, the experiments performed the execution of the decoding algorithm by dealing with the measured data in a post-processing stage, i.e. no real-time decoding was actually performed. However, these experiments represent the state-of-the-art of experimental implementations of surface codes and they are an effort to progress towards the ultimate goal of fault-tolerant quantum computes.